A boy throws a rock straight up into the air. It reaches the highest point of its flight after 2.5 seconds. How fast was the rock giong when it left the boys hand?
Answer by venugopalramana(3286)

THE FORMULA TO BE USED IS
V=U+AT...WHERE
V=FINAL VELOCITY....FT/SEC.
U=INITIAL VELOCITY....FT/SEC.
A = ACCELERATION ..HERE IT IS DUE TO GRAVITY WHICH IS ACTING AGAINST THE MOTION ..ITS VALUE IS -32 FT/SEC.SEC.
T=TIME OF TRAVEL...SEC
WE HAVE U=?
V=0....AS IT REACHED THE HIGHEST POINT AND WILL STOP AND FALL BACK FROM THERE.
T=2.5 SEC.
0=U-32*2.5
U=32*2.5=80 FT/ SEC.
IF YOU ARE IN MKS UNITS THEN A=-9.8 M/SEC.SEC
U=9.8*2.5=24.5 M/SEC